### v1

@@ 5 Object Detection in Autonomous Vehicles
ViKIENet [263] is a multi-modal 3D object detection framework that replaces the dense, noise-prone virtual points of traditional LiDAR–camera fusion with a compact set of semantically enriched Virtual Key Instances (VKIs). Its architecture comprises three tightly integrated modules. The Semantic Key Instance Selection (SKIS) module generates VKIs by combining depth completion with semantic segmentation, converting only object-relevant pixels into 3D points augmented with class scores. These VKIs and raw LiDAR points are voxelized in separate branches, processed by sparse 3D convolutions, and aligned through the Virtual-Instance-to-Real Attention (VIRA) module, which refines VKI features using precise LiDAR depth cues. The Virtual Instance Focused Fusion (VIFF) module then performs two-stage fusion: BEV-level integration via channel and spatial attention to enhance proposal generation, and RoI-level bi-directional cross-attention to combine VKI semantic richness with LiDAR geometric accuracy for final bounding box refinement. This design reduces computation, suppresses depth-induced noise, and preserves discriminative spatial–semantic cues.
@@
PointPainting [283] is a sequential fusion mechanism that decorates every LiDAR point with image semantics before any LiDAR-only detector processes it. An image segmentation network produces per-pixel class scores; LiDAR points are projected into that score map and “painted” by appending the corresponding class-probability vector, yielding an augmented point cloud that remains compatible with arbitrary LiDAR backbones (e.g., PointPillars, VoxelNet/SECOND, PointRCNN). The method’s key property is modularity: fusion is realized as feature augmentation at the point level, enabling independent advancement of the segmentation and detection stacks and practical pipelining for low-latency deployment. SupFusion [272] is a training strategy for LiDAR–camera detectors that introduces auxiliary feature-level supervision from an assistant model trained on densified LiDAR via Polar Sampling. During training, standard 3D/2D backbones and a deep fusion module (stacked MLPs with dynamic fusion blocks) produce fusion features that are optimized not only by the detection loss but also to mimic high-quality features generated by the assistant on enhanced data; this supervision strengthens fusion layers and the upstream encoders without incurring inference overhead. The pipeline thus pairs decision-level supervision with auxiliary alignment of fusion representations, improving robustness of multi-modal feature formation and downstream detection.
@@
Table 22 presents an analysis of 2D–3D fusion detection methods on the paired image–LiDAR NuScenes dataset. We provide a detailed discussion of four selected models: MV2DFusion [288], CLS-3D [264], and ECL3D [289], which achieved the highest results among all the methods in terms of mAP and NDS metrics on the NuScenes dataset; and TransFusion [248], recognized as one of the popular model in the context of object detection. MV2DFusion [288] employs separate 2D and 3D detection backbones, ResNet-50 with a conventional image detector and FSDv2 for LiDAR, to independently produce candidate object proposals for each modality. These proposals are transformed into modality-specific object queries by dedicated query generation modules. A six-layer transformer-based fusion decoder then processes the queries, first modeling relationships within each set via self-attention, and subsequently enabling feature exchange across modalities through cross-attention. A query calibration stage refines the fused representations before final prediction. By focusing on proposal-level fusion, this design maintains the semantic integrity to avoids the pitfalls of dense feature concatenations.
@@
CLS-3D [264] integrates LiDAR and RGB information through a unified backbone in which each modality has its encoder but contributes to a shared feature space for joint reasoning. The image pathway applies DeepLabv3+ to obtain pixel-wise semantic probability maps geometrically aligned to the LiDAR frame. Then, each 3D point is annotated with the semantic probabilities of its corresponding image pixel, creating an enriched point set that combines geometric and class-level cues. This representation is processed by a transformer module equipped with a slot-based reweighting mechanism, allowing the network to emphasise channels that carry the most relevant spatial or temporal information. For final detection, the network employs an I3C-IoU loss considering box overlap, centre displacement, and scale, producing more precise localisation in crowded or visually challenging driving environments.
@@
[289] ECL3D 2025 T-ITS ResNet-50 VoxelNet+ 72.8 75.2 0.252 0.235 0.326 0.184 0.127
@@
[264] CLS-3D 2025 IEEE DeepLabv3+ 3D backbone 75.1 76.8 0.273 0.244 0.302 0.214 0.108
@@
[290] DAL 2024 ECCV ResNet-50 VoxelNet+ 71.5 74.0 0.253 0.239 0.334 0.174 0.120
@@
[288] MV2DFusion-v1 2024 ArXiv ResNet-50 FSDv2 74.5 76.7 0.245 0.229 0.269 0.119 0.115
@@
[288] MV2DFusion-v2 2024 ArXiv ResNet-50 FSDv2 77.9 78.8 0.237 0.226 0.247 0.192 0.119
@@
[291] BEVFusion 2023 ICRA Swin-Tiny VoxelNet 70.2 72.9 0.261 0.239 0.329 0.260 0.134
@@
[292] SparseFusion 2023 ICCV ResNet-50 VoxelNet+ 72.0 73.8 0.258 0.243 0.329 0.265 0.131
@@
[293] CMT 2023 ICCV VOVNet VoxelNet 72.0 74.1 0.279 0.235 0.308 0.259 0.112
@@
[294] FUTR3D 2023 CVPR VoVNet VoxelNet 69.4 72.1 0.284 0.241 0.310 0.300 0.120
@@
[295] FocalFormer3D 2023 ICCV ResNet-50 VoxelNet 72.9 75.0 0.250 0.242 0.328 0.226 0.126
@@
[242] LargeKernel3D-F 2023 CVPR SW-LK SW-LK 71.1 74.2 0.236 0.228 0.298 0.241 0.131
@@
[296] FusionFormer 2023 ArXiv VoV-99 VoxelNet 72.6 75.1 0.267 0.236 0.328 0.226 0.105
@@
[186] UVTR-M 2022 NeurIPS ResNet-101 VoxelNet+ 67.1 71.1 0.306 0.245 0.351 0.225 0.124
@@
[248] TrasFusion 2022 CVPR ResNet-50 VoxelNet 68.9 71.7 0.259 0.243 0.329 0.288 0.127
@@
[298] FusionPainting 2021 ITSC ResNet-50 VoxelNet+ 66.3 70.4 0.256 0.236 0.346 0.274 0.132
@@
[300] PointAugmenting 2021 CVPR DLA-34 VoxelNet+ 66.8 71.1 0.253 0.234 0.345 0.266 0.123
@@
[301] 3D-CVF 2020 ECCV ResNet-18 3D Sparse CNN 67.5 57.8 0.280 0.246 0.367 0.239 0.129
@@
ECL3D [289] follows a dense-to-sparse fusion pipeline with parallel image and LiDAR streams producing BEV features via ResNet-50+FPN with an LSS view transformer and a VoxelNet+SECOND backbone, respectively. The dense stage fuses BEV features to generate a category heatmap, from which top-K proposals are selected. In the sparse stage, dual-modal proposal features are extracted for final regression and classification. Two core enhancements improve accuracy: a dual-scale depth-semantic supervision module in the image branch augments LSS with an auxiliary high-resolution depth prediction branch without increasing inference cost, and an Instance BEV Feature Enhancement (IBFE) module applies deformable cross-attention between LiDAR and image BEVs to align semantic and geometric cues before proposal refinement. TransFusion [248] extracts LiDAR BEV features via a VoxelNet backbone and image features via a ResNet-50 image backbone, then applies a two-stage transformer decoder as the detection head. The first stage decodes sparse, category-aware object queries from LiDAR BEV features into initial bounding boxes. The second stage performs LiDAR–camera fusion using Spatially Modulated Cross-Attention (SMCA), which imposes a locality bias by focusing on relevant image regions for each query. An image-guided query initialization pathway seeds queries using 2D features to help detect small objects under sparse LiDAR sampling. This strategy improves robustness against sensor misalignment and degraded image conditions.
@@ 5.4 LLMs/VLMs-based Methods
LLMs are being recognized as powerful tools for autonomous driving systems, since they facilitate semantic understanding, contextual reasoning and support the extraction of useful information from language data. Their capabilities have been demonstrated across various aspects of autonomous driving, from interpreting sensor data and generating scene summaries to assisting in trajectory planning and decision-making under dynamic traffic conditions. LLMs take a fundamentally different approach from traditional computer vision systems. While conventional CV models depend entirely on what they can ”see” in images, LLMs work with human language. Meanwhile, VLMs are designed to understand and process both images and text together, allowing for more complete and meaningful interpretation of complex driving environments and scenarios. Unlike traditional CV models that only analyze image pixels or LLMs that focus on text, VLMs handle both at the same time. They can describe images, answer questions based on what they ”see,” and retrieve relevant visuals given a text query [302].
@@
[303] 2025 LiDAR-LLM NuScenes-QA Encoder Text, LiDAR Accuracy Reasoning
@@
[304] 2025 BEV-TSR NuScenes Encoder Text, Vision Accuracy Reasoning
@@
[305] 2025 LC-LLM HighD Encoder Text F1, Accuracy Reasoning
@@
[306] 2025 Traj-LLM NuScenes GPT-2, LoRA Text minADE, MR Reasoning
@@
[307] 2024 DriveGPT4 BDD-X Multimodal Text, Video BLEU4 Reasoning
@@
[308] 2024 LLaVA1.5 VQA-v2 Multimodal Text, Video Accuracy, BLEU Reasoning
@@
[309] 2024 GPT-3.5 Generated Multimodal Text Accuracy Captioning
@@
[310] 2023 GAIA-1 Generated Encoder-Decoder Text, Video Cross-entropy Understanding
@@
[311] 2023 GPT-3 Generated Multimodal Text, Vision CLIPort Reasoning
@@
[312] 2023 DriveMLM CARLA Town05 Multimodal LiDAR, Vision F1, BLEU4 Reasoning
@@
[313] 2023 ADAPT Generated Multimodal Text, Video BLEU-1 Reasoning
@@
[314] 2023 GPT-DRIVER NuScenes Multimodal LiDAR, Vision L2 Reasoning
@@
In [303], the authors use LLMs to interpret sparse outdoor 3D LiDAR data. The proposed LiDAR-LLM reformulates 3D scene understanding tasks into a language modeling paradigm, such as captioning, grounding, and question answering. The authors design a three-stage training pipeline involving cross-modal alignment, object-centric perception, and high-level instruction tuning. The Position-Aware Transformer is central to the model, which integrates bird’s-eye view positional embeddings to align 3D LiDAR features with LLM token space. They also construct large-scale synthetic datasets (420K captioning and 280K grounding pairs) using nuScenes and show that LiDAR-LLM outperforms 2D vision-language baselines in multiple 3D tasks, achieving strong results on both their synthetic benchmarks and real-world datasets like NuScenes-QA. This work marks a significant step toward general-purpose, instruction-following LLMs for 3D scene understanding in AVs. The method proposed in the [306] proposes a trajectory prediction in autonomous driving using pre-trained Large Language Models without explicit prompt engineering. The model captures high-level semantic interactions and scene context by encoding spatial-temporal features of agents and traffic scenes into LLM formats. It incorporates a novel lane-aware probabilistic learning module, powered by the Mamba architecture, to simulate human-like attention to lane structures. Moreover, a multi-modal Laplace decoder generates diverse and scene-compliant future trajectories. The results show that this method outperforms state-of-the-art baselines across key metrics, including in few-shot settings. It highlights its robustness, generalizability, and the untapped potential of LLMs for structured prediction in complex driving environments. [78] presents a method that integrates VLMs into end-to-end autonomous driving by modeling human-like reasoning through Graph Visual Question Answering. This method captures multistep, multiobject dependencies by structuring perception, prediction, and planning tasks as interconnected Question Answering pairs in a graph. Moreover, the authors propose two datasets, DriveLM-nuScenes and DriveLM-CARLA, and a baseline model, DriveLM-Agent, which combines graph prompting with trajectory tokenization to repurpose general VLMs for driving tasks. Results demonstrate that this method performs well with driving-specific models and excels in zero-shot generalization, especially with unseen sensor configurations. This work highlights Graph VQA as a promising approach for enabling explainable, generalizable, and interactive AV system.
@@
[315] 2025 DynRsl-VLM ViT+ResNet+Transformer NuInstruct Text, Image Accuracy, BLEU Reasoning
@@
[316] 2025 AV-VLM GPT-4V Generated Text, Vision, LiDAR Accuracy, Latency Reasoning
@@
[317] 2025 VLC Fusion BEVFusion + CLIP Waymo, ATR Text, Vision, LiDAR IR, mAP Detection
@@
[318] 2025 COOOL LLaMA-2 + CLIP COOOLER Video, Text BESM, SAM Captioning
@@
[319] 2025 NuPrompt Transformer NuPrompt Text, Vision AMOTA, ADE Prediction
@@
[320] 2024 Llama-Adapter Multimodal-7B Generated Text, Vision Accuracy, BLEU Reasoning
@@
[321] 2024 VLM-RL LLaMA+CLIP Generated Vision, Text AS, RC Reward Design
@@
[322] 2024 VLM-AD LLaMA-2 NuScenes Vision, Text L2 Error, Collision Rate Reasoning
@@
[323] 2024 EM-VLM4AD Multiple DriveLM Image, Text BLEU-4, METEOR Reasoning
@@
[78] 2023 DriveLM-Agent BLIP-2 NuScenes Text, Image Accuracy Reasoning
@@
[324] 2023 BLIP-2 LLaMA-2-7B COCO, VQAv2 Text, Vision Accuracy, BLEU Captioning
@@
[325] 2024 BEV-CLIP CLIP+Llama2 NuScenes Text, BEV Accuracy, Captioning
@@
[326] 2024 Talk2BEV BEVFormer+GPT-2 NuScenes Text, Image, LiDAR Accuracy, IoU Reasoning
@@
[327] 2023 MSSG GPT-2 NuScenes LiDAR, Text BEV AP, 3D AP Reasoning
@@
[328] 2023 RMOT Swin-L+BERT KITTI Vision HOTA, DetA, Tracking
@@
[329] 2023 CLIP2Scene CLIP+PointNet++ NuScenes, ScanNet Text, Vision mIoU, Accuracy Reasoning
@@
[330] 2023 CLIP ViT+Transformer Waymo Text, LiDAR 3D AP Detection, Reasoning

### v2

@@ 5 Object Detection in Autonomous Vehicles
ViKIENet [yu2025vikienet] is a multi-modal 3D object detection framework that replaces the dense, noise-prone virtual points of traditional LiDAR–camera fusion with a compact set of semantically enriched Virtual Key Instances (VKIs). Its architecture comprises three tightly integrated modules. The Semantic Key Instance Selection (SKIS) module generates VKIs by combining depth completion with semantic segmentation, converting only object-relevant pixels into 3D points augmented with class scores. These VKIs and raw LiDAR points are voxelized in separate branches, processed by sparse 3D convolutions, and aligned through the Virtual-Instance-to-Real Attention (VIRA) module, which refines VKI features using precise LiDAR depth cues. The Virtual Instance Focused Fusion (VIFF) module then performs two-stage fusion: BEV-level integration via channel and spatial attention to enhance proposal generation, and RoI-level bi-directional cross-attention to combine VKI semantic richness with LiDAR geometric accuracy for final bounding box refinement. This design reduces computation, suppresses depth-induced noise, and preserves discriminative spatial–semantic cues.
@@
PointPainting [vora2020pointpainting] is a sequential fusion mechanism that decorates every LiDAR point with image semantics before any LiDAR-only detector processes it. An image segmentation network produces per-pixel class scores; LiDAR points are projected into that score map and “painted” by appending the corresponding class-probability vector, yielding an augmented point cloud that remains compatible with arbitrary LiDAR backbones (e.g., PointPillars, VoxelNet/SECOND, PointRCNN). The method’s key property is modularity: fusion is realized as feature augmentation at the point level, enabling independent advancement of the segmentation and detection stacks and practical pipelining for low-latency deployment. SupFusion [qin2023supfusion] is a training strategy for LiDAR–camera detectors that introduces auxiliary feature-level supervision from an assistant model trained on densified LiDAR via Polar Sampling. During training, standard 3D/2D backbones and a deep fusion module (stacked MLPs with dynamic fusion blocks) produce fusion features that are optimized not only by the detection loss but also to mimic high-quality features generated by the assistant on enhanced data; this supervision strengthens fusion layers and the upstream encoders without incurring inference overhead. The pipeline thus pairs decision-level supervision with auxiliary alignment of fusion representations, improving robustness of multi-modal feature formation and downstream detection.
@@
Table 22 presents an analysis of 2D–3D fusion detection methods on the paired image–LiDAR NuScenes dataset. We provide a detailed discussion of four selected models: MV2DFusion [wang2024mv2dfusion], CLS-3D [mushtaq2025cls], and ECL3D [chen2025multi], which achieved the highest results among all the methods in terms of mAP and NDS metrics on the NuScenes dataset; and TransFusion [bai2022transfusion], recognized as one of the popular model in the context of object detection. MV2DFusion [wang2024mv2dfusion] employs separate 2D and 3D detection backbones, ResNet-50 with a conventional image detector and FSDv2 for LiDAR, to independently produce candidate object proposals for each modality. These proposals are transformed into modality-specific object queries by dedicated query generation modules. A six-layer transformer-based fusion decoder then processes the queries, first modeling relationships within each set via self-attention, and subsequently enabling feature exchange across modalities through cross-attention. A query calibration stage refines the fused representations before final prediction. By focusing on proposal-level fusion, this design maintains the semantic integrity to avoids the pitfalls of dense feature concatenations.
@@
CLS-3D [mushtaq2025cls] integrates LiDAR and RGB information through a unified backbone in which each modality has its encoder but contributes to a shared feature space for joint reasoning. The image pathway applies DeepLabv3+ to obtain pixel-wise semantic probability maps geometrically aligned to the LiDAR frame. Then, each 3D point is annotated with the semantic probabilities of its corresponding image pixel, creating an enriched point set that combines geometric and class-level cues. This representation is processed by a transformer module equipped with a slot-based reweighting mechanism, allowing the network to emphasise channels that carry the most relevant spatial or temporal information. For final detection, the network employs an I3C-IoU loss considering box overlap, centre displacement, and scale, producing more precise localisation in crowded or visually challenging driving environments.
@@
[chen2025multi] ECL3D 2025 T-ITS ResNet-50 VoxelNet+ 72.8 75.2 0.252 0.235 0.326 0.184 0.127
@@
[mushtaq2025cls] CLS-3D 2025 IEEE DeepLabv3+ 3D backbone 75.1 76.8 0.273 0.244 0.302 0.214 0.108
@@
[huang2024detecting] DAL 2024 ECCV ResNet-50 VoxelNet+ 71.5 74.0 0.253 0.239 0.334 0.174 0.120
@@
[wang2024mv2dfusion] MV2DFusion-v1 2024 ArXiv ResNet-50 FSDv2 74.5 76.7 0.245 0.229 0.269 0.119 0.115
@@
[wang2024mv2dfusion] MV2DFusion-v2 2024 ArXiv ResNet-50 FSDv2 77.9 78.8 0.237 0.226 0.247 0.192 0.119
@@
[liu2022bevfusion] BEVFusion 2023 ICRA Swin-Tiny VoxelNet 70.2 72.9 0.261 0.239 0.329 0.260 0.134
@@
[xie2023sparsefusion] SparseFusion 2023 ICCV ResNet-50 VoxelNet+ 72.0 73.8 0.258 0.243 0.329 0.265 0.131
@@
[yan2023cross] CMT 2023 ICCV VOVNet VoxelNet 72.0 74.1 0.279 0.235 0.308 0.259 0.112
@@
[chen2023futr3d] FUTR3D 2023 CVPR VoVNet VoxelNet 69.4 72.1 0.284 0.241 0.310 0.300 0.120
@@
[chen2023focalformer3d] FocalFormer3D 2023 ICCV ResNet-50 VoxelNet 72.9 75.0 0.250 0.242 0.328 0.226 0.126
@@
[chen2023largekernel3d] LargeKernel3D-F 2023 CVPR SW-LK SW-LK 71.1 74.2 0.236 0.228 0.298 0.241 0.131
@@
[hu2023fusionformer] FusionFormer 2023 ArXiv VoV-99 VoxelNet 72.6 75.1 0.267 0.236 0.328 0.226 0.105
@@
[li2022unifying] UVTR-M 2022 NeurIPS ResNet-101 VoxelNet+ 67.1 71.1 0.306 0.245 0.351 0.225 0.124
@@
[bai2022transfusion] TrasFusion 2022 CVPR ResNet-50 VoxelNet 68.9 71.7 0.259 0.243 0.329 0.288 0.127
@@
[xu2021fusionpainting] FusionPainting 2021 ITSC ResNet-50 VoxelNet+ 66.3 70.4 0.256 0.236 0.346 0.274 0.132
@@
[wang2021pointaugmenting] PointAugmenting 2021 CVPR DLA-34 VoxelNet+ 66.8 71.1 0.253 0.234 0.345 0.266 0.123
@@
[yoo20203d] 3D-CVF 2020 ECCV ResNet-18 3D Sparse CNN 67.5 57.8 0.280 0.246 0.367 0.239 0.129
@@
ECL3D [chen2025multi] follows a dense-to-sparse fusion pipeline with parallel image and LiDAR streams producing BEV features via ResNet-50+FPN with an LSS view transformer and a VoxelNet+SECOND backbone, respectively. The dense stage fuses BEV features to generate a category heatmap, from which top-K proposals are selected. In the sparse stage, dual-modal proposal features are extracted for final regression and classification. Two core enhancements improve accuracy: a dual-scale depth-semantic supervision module in the image branch augments LSS with an auxiliary high-resolution depth prediction branch without increasing inference cost, and an Instance BEV Feature Enhancement (IBFE) module applies deformable cross-attention between LiDAR and image BEVs to align semantic and geometric cues before proposal refinement. TransFusion [bai2022transfusion] extracts LiDAR BEV features via a VoxelNet backbone and image features via a ResNet-50 image backbone, then applies a two-stage transformer decoder as the detection head. The first stage decodes sparse, category-aware object queries from LiDAR BEV features into initial bounding boxes. The second stage performs LiDAR–camera fusion using Spatially Modulated Cross-Attention (SMCA), which imposes a locality bias by focusing on relevant image regions for each query. An image-guided query initialization pathway seeds queries using 2D features to help detect small objects under sparse LiDAR sampling. This strategy improves robustness against sensor misalignment and degraded image conditions.
@@ 5.4 LLMs/VLMs-based Methods
LLMs are being recognized as powerful tools for autonomous driving systems, since they facilitate semantic understanding, contextual reasoning and support the extraction of useful information from language data. Their capabilities have been demonstrated across various aspects of autonomous driving, from interpreting sensor data and generating scene summaries to assisting in trajectory planning and decision-making under dynamic traffic conditions. LLMs take a fundamentally different approach from traditional computer vision systems. While conventional CV models depend entirely on what they can ”see” in images, LLMs work with human language. Meanwhile, VLMs are designed to understand and process both images and text together, allowing for more complete and meaningful interpretation of complex driving environments and scenarios. Unlike traditional CV models that only analyze image pixels or LLMs that focus on text, VLMs handle both at the same time. They can describe images, answer questions based on what they ”see,” and retrieve relevant visuals given a text query [vahidi2025shadow].
@@
[yang2025lidar] 2025 LiDAR-LLM NuScenes-QA Encoder Text, LiDAR Accuracy Reasoning
@@
[tang2025bev] 2025 BEV-TSR NuScenes Encoder Text, Vision Accuracy Reasoning
@@
[peng2025lc] 2025 LC-LLM HighD Encoder Text F1, Accuracy Reasoning
@@
[lan2024traj] 2025 Traj-LLM NuScenes GPT-2, LoRA Text minADE, MR Reasoning
@@
[xu2024drivegpt4] 2024 DriveGPT4 BDD-X Multimodal Text, Video BLEU4 Reasoning
@@
[liu2024improved] 2024 LLaVA1.5 VQA-v2 Multimodal Text, Video Accuracy, BLEU Reasoning
@@
[chen2024driving] 2024 GPT-3.5 Generated Multimodal Text Accuracy Captioning
@@
[hu2023gaia] 2023 GAIA-1 Generated Encoder-Decoder Text, Video Cross-entropy Understanding
@@
[elhafsi2023semantic] 2023 GPT-3 Generated Multimodal Text, Vision CLIPort Reasoning
@@
[wang2023drivemlm] 2023 DriveMLM CARLA Town05 Multimodal LiDAR, Vision F1, BLEU4 Reasoning
@@
[jin2023adapt] 2023 ADAPT Generated Multimodal Text, Video BLEU-1 Reasoning
@@
[mao2023gpt] 2023 GPT-DRIVER NuScenes Multimodal LiDAR, Vision L2 Reasoning
@@
In [yang2025lidar], the authors use LLMs to interpret sparse outdoor 3D LiDAR data. The proposed LiDAR-LLM reformulates 3D scene understanding tasks into a language modeling paradigm, such as captioning, grounding, and question answering. The authors design a three-stage training pipeline involving cross-modal alignment, object-centric perception, and high-level instruction tuning. The Position-Aware Transformer is central to the model, which integrates bird’s-eye view positional embeddings to align 3D LiDAR features with LLM token space. They also construct large-scale synthetic datasets (420K captioning and 280K grounding pairs) using nuScenes and show that LiDAR-LLM outperforms 2D vision-language baselines in multiple 3D tasks, achieving strong results on both their synthetic benchmarks and real-world datasets like NuScenes-QA. This work marks a significant step toward general-purpose, instruction-following LLMs for 3D scene understanding in AVs. The method proposed in the [lan2024traj] proposes a trajectory prediction in autonomous driving using pre-trained Large Language Models without explicit prompt engineering. The model captures high-level semantic interactions and scene context by encoding spatial-temporal features of agents and traffic scenes into LLM formats. It incorporates a novel lane-aware probabilistic learning module, powered by the Mamba architecture, to simulate human-like attention to lane structures. Moreover, a multi-modal Laplace decoder generates diverse and scene-compliant future trajectories. The results show that this method outperforms state-of-the-art baselines across key metrics, including in few-shot settings. It highlights its robustness, generalizability, and the untapped potential of LLMs for structured prediction in complex driving environments. [sima2024drivelm] presents a method that integrates VLMs into end-to-end autonomous driving by modeling human-like reasoning through Graph Visual Question Answering. This method captures multistep, multiobject dependencies by structuring perception, prediction, and planning tasks as interconnected Question Answering pairs in a graph. Moreover, the authors propose two datasets, DriveLM-nuScenes and DriveLM-CARLA, and a baseline model, DriveLM-Agent, which combines graph prompting with trajectory tokenization to repurpose general VLMs for driving tasks. Results demonstrate that this method performs well with driving-specific models and excels in zero-shot generalization, especially with unseen sensor configurations. This work highlights Graph VQA as a promising approach for enabling explainable, generalizable, and interactive AV system.
@@
[zhou2025dynrsl] 2025 DynRsl-VLM ViT+ResNet+Transformer NuInstruct Text, Image Accuracy, BLEU Reasoning
@@
[zhou2025hierarchical] 2025 AV-VLM GPT-4V Generated Text, Vision, LiDAR Accuracy, Latency Reasoning
@@
[taparia2025vlc] 2025 VLC Fusion BEVFusion + CLIP Waymo, ATR Text, Vision, LiDAR IR, mAP Detection
@@
[shriram2025towards] 2025 COOOL LLaMA-2 + CLIP COOOLER Video, Text BESM, SAM Captioning
@@
[wu2025language] 2025 NuPrompt Transformer NuPrompt Text, Vision AMOTA, ADE Prediction
@@
[he2024integrating] 2024 Llama-Adapter Multimodal-7B Generated Text, Vision Accuracy, BLEU Reasoning
@@
[huang2024vlm] 2024 VLM-RL LLaMA+CLIP Generated Vision, Text AS, RC Reward Design
@@
[xu2024vlm] 2024 VLM-AD LLaMA-2 NuScenes Vision, Text L2 Error, Collision Rate Reasoning
@@
[gopalkrishnan2024multi] 2024 EM-VLM4AD Multiple DriveLM Image, Text BLEU-4, METEOR Reasoning
@@
[sima2024drivelm] 2023 DriveLM-Agent BLIP-2 NuScenes Text, Image Accuracy Reasoning
@@
[li2023blip] 2023 BLIP-2 LLaMA-2-7B COCO, VQAv2 Text, Vision Accuracy, BLEU Captioning
@@
[jia2024bev] 2024 BEV-CLIP CLIP+Llama2 NuScenes Text, BEV Accuracy, Captioning
@@
[choudhary2024talk2bev] 2024 Talk2BEV BEVFormer+GPT-2 NuScenes Text, Image, LiDAR Accuracy, IoU Reasoning
@@
[cheng2023language] 2023 MSSG GPT-2 NuScenes LiDAR, Text BEV AP, 3D AP Reasoning
@@
[wu2023referring] 2023 RMOT Swin-L+BERT KITTI Vision HOTA, DetA, Tracking
@@
[chen2023clip2scene] 2023 CLIP2Scene CLIP+PointNet++ NuScenes, ScanNet Text, Vision mIoU, Accuracy Reasoning
@@
[najibi2023unsupervised] 2023 CLIP ViT+Transformer Waymo Text, LiDAR 3D AP Detection, Reasoning